# ViSurf: Visual Supervised-and-Reinforcement Fine-Tuning for Large Vision-and-Language Models

### v2

@@ 4.5 Comparison with State-of-The-Arts
We compare ViSurf against state-of-the-art (SoTA) models on two visual perception tasks: gRefCOCO and ReasonSeg. We compare LISA [lai2024lisa], GSVA [xia2024gsva], SAM4MLLM [chen2024sam4mllm], SegZero [liu2025seg], VisionReasoner [liu2025visionreasoner]. As shown in LABEL:{tab:compare_sotas}, ViSurf achieves the state-of-the-arts performance on ReasonSeg and gRefCOCO.
@@ 4.6 Qualitative Results
Qualitative results on various tasks are presented in Figure 7. The proposed method, ViSurf, successfully localizes normal objects, identifies anomalies, detects non-object cases, and solves mathematical problems.
@@ 5 Conclusion
We propose ViSurf, a single-stage post-training paradigm that integrates the benefits of both SFT and RLVR, motivated by a theoretical analysis of their objectives and gradients. The practical implementation of ViSurf involves interleaving ground-truth labels with model-generated rollouts, augmented by three reward control strategies to ensure training stability. Experimental results across diverse benchmarks demonstrate that ViSurf outperforms SFT, RLVR, and a sequential SFT → RLVR pipeline. The subsequent in-depth analysis provides further insights and corroborates the theoretical analysis.
[…]
Supplementary Material
@@ Appendix C Non Object Training Data
The training data for gRefCOCO was adapted from the VisionReasoner training data [liu2025visionreasoner], from which we utilized 7k referring expression samples. To enhance the model’s ability to handle cases where no target object is present, we augmented this dataset with 200 non-object examples. These negative samples were generated by providing a question that is unanswerable given the image content and training the model to output an empty list (<answer>[]</answer>).
@@ Appendix D Illustration of Prompt Design
Table 9 shows prompt with detailed format instruction, where we provide desired output format for model. Table 10 shows prompt without detailed format instruction, where we simply write ‘answer here’ between answer tags.
Table 9: Prompt with detailed format instruction. We provide detailed format instructions between answer tags.
Prompt with Detailed Instruction ”Please find “{Question}” with bboxs and points.” ”Compare the difference between object(s) and find the most closely matched object(s).” ”Output the thinking process in <<think>> <</think>> and final answer in <<answer>> <</answer>> tags.” ”Output the bbox(es) and point(s) inside the interested object(s) in JSON format.”
i.e. <think> thinking process here </think>
<answer>[{"bbox_2d": [10, 100, 200, 210], "point_2d" [30, 110]},
{"bbox_2d": [225, 296, 706, 786], "point_2d": [302, 410]}]</answer>
Table 10: Prompt without Detailed Instruction. We do not provide detailed format instructions between answer tags.
Prompt without Detailed Instruction ”Please find “{Question}” with bboxs and points.” ”Compare the difference between object(s) and find the most closely matched object(s).” ”Output the thinking process in <<think>> <</think>> and final answer in <<answer>> <</answer>> tags.” ”Output the bbox(es) and point(s) inside the interested object(s) in JSON format.”
i.e. <think> thinking process here </think>
<answer> answer here </answer>

### v3

@@ 4.5 Comparison with State-of-The-Arts
We compare ViSurf against state-of-the-art (SoTA) models on two visual perception tasks: gRefCOCO and ReasonSeg. We compare LISA (Lai et al., 2024), GSVA (Xia et al., 2024), SAM4MLLM (Chen et al., 2024a), SegZero (Liu et al., 2025b), VisionReasoner (Liu et al., 2025c). As shown in LABEL:{tab:compare_sotas}, ViSurf achieves the SoTA performance.
@@ 4.6 Qualitative Results
Qualitative results on various tasks are presented in Figure 7. The results demonstrate that models trained with ViSurf successfully solves multiple visual perception tasks.
@@ 5 Conclusion
We present ViSurf, a unified post-training paradigm designed to bridge the gap between Supervised Fine-Tuning (SFT) and Reinforcement Learning from Verifiable Rewards (RLVR). Driven by a rigorous theoretical analysis of their respective objectives and gradient dynamics, ViSurf integrates the strengths of both approaches into a single-stage framework. Our implementation strategically interleaves ground-truth labels with model-generated rollouts, utilizing three novel reward control strategies to maintain training stability. Empirical evaluations across diverse benchmarks demonstrate that ViSurf consistently outperforms standalone SFT, RLVR, and the traditional sequential SFT $\rightarrow$ RLVR pipeline. Our in-depth analysis provides deep insights into the framework’s behavior, effectively corroborating our theoretical foundations.
Impact Statements
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix C Non Object Training Data
The training data for gRefCOCO was adapted from the VisionReasoner training data (Liu et al., 2025c), from which we utilized 7k referring expression samples. To enhance the model’s ability to handle cases where no target object is present, we augmented this dataset with 200 non-object examples. These negative samples were generated by providing a question that is unanswerable given the image content and training the model to output an empty list (<answer>[]</answer>).
@@ Appendix D Illustration of Prompt Design
Prompt D shows prompt with detailed format instruction, where we provide desired output format for model. Prompt D shows prompt without detailed format instruction, where we simply write ‘answer here’ between answer tags.
Prompt 1: with detailed format instruction “Please find ‘{{Question}}}’ with bboxs and points.” “Compare the difference between object(s) and find the most closely matched object(s).” “Output the thinking process in <think> </think> and final answer in <answer> </answer> tags.” “Output the bbox(es) and point(s) inside the interested object(s) in JSON format.”
i.e. <think>
thinking process here
</think>
<answer>
[{"bbox_2d": [10, 100, 200, 210],
"point_2d": [30, 110]},
{"bbox_2d": [225, 296, 706, 786],
"point_2d": [302, 410]}]
</answer>
Prompt 2: without detailed format instruction “Please find ‘{{Question}}}’ with bboxs and points.” “Compare the difference between object(s) and find the most closely matched object(s).” “Output the thinking process in <think> </think> and final answer in <answer> </answer> tags.” “Output the bbox(es) and point(s) inside the interested object(s) in JSON format.”
i.e. <think>
thinking process here
</think>
<answer>
answer here
</answer>
Appendix E Detailed Experimental Settings
Non-Object Segmentation. The gRefCOCO (Liu et al., 2023a) includes queries that do not contain corresponding objects. The evaluation metrics are gIoU and N-Acc. We use Multi-objects-7K (Liu et al., 2025c) plus with 200 non-object data for training.
Reasoning Segmentation. The ReasonSeg (Lai et al., 2024) includes test samples that need reasoning for correct segmentation. It has 200 validation images and 779 test images. The evaluation metric is gIoU. We use Multi-objects-7K (Liu et al., 2025c) proposed in VisionReasoner (Liu et al., 2025c) for training.
GUI Grounding. The OmniACT (Kapoor et al., 2024) is a GUI grounding task for Desktop and Web. We randomly collect 6,101 samples in training split and verify on the test split. The accuracy is calculated as whether the predict point correctly locates inside the interest region.
Anomaly Detection. The RealIAD (Wang et al., 2024a) includes real-world, multi-view industrial anomaly. We derive 3,292 training samples and 2,736 test samples, ensuring the two sets are disjoint. We calculated the ROC_AUC.
Medical Image: Skin. The task one of ISIC2018 (Codella et al., 2019; International Skin Imaging Collaboration (ISIC), 2018) is lesion segmentation. It includes 2,594 training samples and 1,000 test samples. We measure the bbox_acc metric, which computes the ratio of predicted bboxes whose IoU with the ground truth exceeds 0.5.
MathVista. The MathVista-testmini (Lu et al., 2023) includes 1,000 diverse mathematical and visual tasks. We gather around 10k training data from WeMath (Qiao et al., 2024), MathVision (Wang et al., 2024b), Polymath (Gupta et al., 2024), SceMQA (Liang et al., 2024), Geometry3K (Lu et al., 2021).